# Absence of twisting for non-trivial discrete torsion

### v1

@@ 2. Partition function of the $n$-torus
Let $\mathcal{T}$ be a $d$-dimensional QFT (or CFT) with finite global symmetry group $G$, and let $\omega]\in H^{d}(G,U(1))$ encode a topological $G$-action (discrete torsion). For a closed oriented $d$-manifold $M$, a flat $G$-bundle AA on $M$ is equivalently a homomorphism $f_{A}:\pi_{1}(M)\to G$, or a map $f_{A}:M\to BG$ up to homotopy. The Dijkgraaf–Witten topological gauge theory with action $\omega$ has partition function
@@ 5.1. Untwisted groups
It has been unknown whether untwisted groups, that is, groups $G$ with $\operatorname{Utw}^{2}(G)=H^{2}(G,U(1))\neq 1$ exist. We have found first known cases of groups $G$ with this property. In GAP [9] notation for small groups, the smallest such group is $[64,182]$ (182182nd group in the library of all groups of order $64$). This is one of the eight groups of order 64 with non-trivial Bogomolov multiplier. There are two groups of order $243$ that are untwisted, namely $[243,29]$ and $[243,30]$. Out of $2328$ groups of order $128$, there are precisely 0 groups that are untwisted.
@@ 6. Conclusions
There are several natural directions for further work. First, it would be valuable to understand $\operatorname{Utw}^{n}(G)$ more conceptually in terms of the homotopy type of the commuting-tuple space and the induced maps $T^{n}\to BG$, potentially relating $H_{0n}(G,\mathbb{Z})$ to a subquotient of $H_{n}(BG)$ detected by maps from tori. Second, extending cocycle-level functionality beyond degree $3$ would enable direct constructions of representative cocycles realizing nontrivial classes in $H^{n}(G,U(1))$ and a more detailed comparison with DW partition functions on manifolds beyond the torus. Finally, on the orbifold/CFT side, it would be interesting to clarify which observables (beyond torus partition functions) can distinguish classes in $H^{n}(G,U(1))$ that are torus-invisible, and whether $\operatorname{Utw}^{n}(G)$ governs a broader notion of “perturbative invisibility” in orbifold theories.

### v2

@@ 2. Partition function of the $n$-torus
Let $\mathcal{T}$ be a $d$-dimensional QFT (or CFT) with finite global symmetry group $G$, and let $\omega]\in H^{d}(G,U(1))$ encode a topological $G$-action (discrete torsion). For a closed oriented $d$-manifold $M$, a flat $G$-bundle AA on $M$ is equivalently a homomorphism $f_{A}:\pi_{1}(M)\to G$, or a map $f_{A}:M\to BG$ up to homotopy. The Dijkgraaf–Witten topological gauge theory [7] with action $\omega$ has partition function
@@ 3. Untwisted discrete torsion and unramified Brauer groups
It is known [16] that if $G_{1}$ and $G_{2}$ are finite groups, then $\operatorname{Utw}^{2}(G_{1}\times G_{2})\cong\operatorname{Utw}^{2}(G_{1})\times\operatorname{Utw}^{2}(G_{2})$. There is a well established fact [17, p. 46] that $H^{2}(G_{1}\times G_{2},U(1))\cong H^{2}(G_{1},U(1))\times H^{2}(G_{2},U(1))\times(G_{1}\otimes G_{2})$, where $G_{1}\otimes G_{2}$ is the tensor product of the largest abelian quotients of $G_{1}$ and $G_{2}$. We conclude the following:
Proposition 3.2.
Let $G_{1}$ and $G_{2}$ be untwisted groups of coprime orders. Then $G_{1}\times G_{2}$ is also untwisted.
@@ 5.1. Untwisted groups
We list some cases of untwisted groups, that is, groups $G$ with $\operatorname{Utw}^{2}(G)=H^{2}(G,U(1))\neq 1$. In GAP [9] notation for small groups, the smallest such group is $[64,182]$ (182182nd group in the library of all groups of order $64$). This is one of the eight groups of order 64 with non-trivial Bogomolov multiplier. There are two groups of order $3^{5}$ that are untwisted, namely $[243,29]$ and $[243,30]$. There are 5 untwisted groups of order $5^{5}$. These are $[3125,34]$, $[3125,35]$, $[3125,36]$, $[3125,37]$, $[3125,38]$. Similarly, the groups $[16807,38]$, $[16807,39]$, $[16807,40]$, $[16807,41]$ and $[16807,42]$ are the utwisted groups of order $7^{5}$. We collect some statistics on other untwisted groups in Table 2.
Order Nr all Nr untwisted
$2^{7}$ 2328 16
$2^{6}\cdot 3$ 1543 6
$2^{6}\cdot 5$ 1640 6
$2^{6}\cdot 7$ 1396 6
$2^{6}\cdot 3^{2}$ 8681 26
$3^{6}$ 504 14
$3^{6}\cdot 2$ 1798 46
$3^{7}$ 9310 153
$5^{6}$ 684 38
Table 2. Untwisted groups.